Find the second-order partial derivative. Find fxy when f(x,y) = 8x®y - 7y2 + 2x. O A. 24x? B. 48xy O C. - 14 OD. -28
Answer #1

Poxy, and order partial derivative fcany)= 8 x ²4 - 74² + 20 fx = 0 l (82²4-7 y2 + 2x) = 240²y + 2 faya [2400²y+2) og 2 = (24